OpenStudy (anonymous):

find the points of intersection of the curves R= 2sina and R= 2sin2a a:angel

OpenStudy (anonymous):

Well, we know that \[ sin(2x) = 2sin(x)cos(x) \] So now to find intersections we can just: \[ 2sin(a) = 2sin(2a) \\ 2sin(a) = 2 * 2sin(a)cos(a) \] Now, we can divide both sides by 2sin(a) as long as a =/= 180K (because sin(180K) = 0 and we can't divide by zero..) \[ \frac{2sin(a)}{2sin(a)} = \frac{2 \cdot 2sin(a)cos(a)}{2sin(a)} \\ 1 = 2cos(a)\\ cos(a) = 0.5\\ a = +-60^\circ + 360^\circ K \] If a = 180K It's easy to see that the following is true: \[ 2sin(a) = 2sin(2a) \\ 2sin(0) = 2sin(0) \\ 0 = 0 \] So, those are the values we got: \[ a = 180^\circ K \\ \text{OR} \\ a = \pm60^\circ + 360^\circ K \] Hope that helps =]